OpenStudy (anonymous):

Find the area of the circle with the given radius or diameter. Use = 3.14. d = 10 A = a0 sq. units

OpenStudy (anonymous):

Areaof circle = pi r ^2 = 3.14*5^2

OpenStudy (anonymous):

A= D/2(3.14) A=10/2(3.14) A=5(3.14) A=15.70

OpenStudy (anonymous):

Thank you